# Subdiffusive transport in the Fredkin dynamical universality class

### v4

@@ B.2 Source of charge-current correlations
In the picture of the matched parentheses charge, it becomes clear that the source of all Fredkin dynamics is the movement of $q^{1}$ charges, which are matched parentheses of the form (). When a $q^{1}$ charge moves, it necessarily disrupts the end parenthesis of a larger $q^{\ell}$ charge. This can be pictured as the $q^{1}$ moving past one of the matched end parentheses of the $q^{\ell}$ charge, which we will call moving “inside” or “outside” of the $q^{\ell}$ charge. The movement of the $q^{1}$ charge triggers the process $q^{\ell}_{n}\rightarrow q^{\ell\pm 1}_{n\pm 1}$.
@@
Since the only source of dynamics from the Fredkin Hamiltonian involve locally moving $q^{1}$ charges, all nontrivial correlations between a charge sitting on link 0 and a current on site $n$ must come from the coincidence of two scenarios: first, that the charge sitting on link 0 corresponds to matched parentheses separated by ±$\ell\pm 1$, and second, that a charge $q_{\pm(\ell\mp 1)}^{1}$ disturbs the end parenthesis of the $q_{0}^{\ell\pm 1}$ charge. This statement has an intuitive interpretation– a current may only be correlated with a $q^{\ell}$ charge if a $q^{1}$ charge either moves inside or outside a $q^{\ell}$ charge, which consequently changes the separation of the charge $q^{\ell}\rightarrow q^{\ell\pm 1}$. In order to move inside or outside of the $q^{\ell}$ charge at a given timestep, since there are no non-local $q^{1}$ hops encoded in the Fredkin Hamiltonian, there must be a $q^{1}$ charge located directly adjacent to one of the matched end parentheses of the $q^{\ell}$ charge immediately before the hop. Therefore, the task of understanding the static charge-current correlations in Fredkin reduces to finding the probability of these configurations. Exact expressions for these probabilities are derived in the following section.
@@ B.3 Probabilities
First, we focus on the process $()(\dots)^{\ell}\rightarrow(()\dots)^{\ell+1}$, in which $(\dots)^{\ell}$ pictorially represents a $q^{\ell}$ charge. We denote the probability of finding a configuration of the form $()(\dots)^{\ell}$ or $(\dots)^{\ell}()$ conditional on having a $(\dots)^{\ell}$ as $p_{\rm out}(\ell)$. From direct counting, we find that $p_{\rm out}(\ell)=7/16$ in the thermodynamic limit since the system is locally at infinite temperature. For a system of $L$ sites in the $S^{z}_{\rm tot}=0$ ensemble, however, the expression for $p_{\rm out}(\ell,L)$ becomes dependent on both $\ell$ and $L$ due to additional correlations that arise from finite-size effects.
Since the Dyck word corresponding to any $q^{\ell}$ charge contains equal numbers of up and down spins, the $L-2\ell$ sites in the system that are not part of the $q^{\ell}$ charge must also contain the same number of up and down spins when the system is at $S^{z}_{\rm tot}=0$. There are $\binom{L-2\ell}{(L-2\ell)/2}$ possible states for these $L-2\ell$ sites. Consider a Dyck word of semilength $\ell+1$ of the form $()(\dots)^{\ell}$; since $()(\dots)^{\ell}$ also contains only matched sets of parentheses, the $L-(\ell+2)$ sites in the system that are not part of $q^{\ell}$ must contain the same number of up and down spins. Therefore, there are $\binom{L-2\ell-2}{(L-2\ell-2)/2}$ possible configurations for these $L-2\ell-2$ sites. Finally, taking care to avoid double-counting the $\binom{L-2\ell-4}{(L-2\ell-4)/2}$ configurations of the form $()(\dots)^{\ell}()$, we obtain the following expression for $p_{\rm out}(\ell,L)$:
@@
Next, we focus on the process $(()\dots)^{\ell}\rightarrow()(\dots)^{\ell-1}$. We denote the probability of finding a configuration of the form $(()\dots)^{\ell}$ conditional on having a $(\dots)^{\ell}$ as $p_{\rm in}(\ell)$. Unlike when a $q^{1}$ charge sits directly outside of a $q^{\ell}$ charge, a $q^{1}$ sitting directly inside of a $q^{\ell}$ charge is subject to strong correlations from being embedded within a Dyck word. In total, there are $C_{\ell-1}$ possible Dyck words corresponding to a $q^{\ell}$ charge, where $C_{\ell}$ are Catalan numbers. Recall that this was obtained from there fact that of the $C_{\ell}$ Dyck word of semilength $\ell$, only those of the form $(\dots)^{\ell}$ correspond to $q^{\ell}$ charges; therefore, the $\dots$ in $(\dots)^{\ell}$ must also be a Dyck word of semilength $\ell-1$, of which there are $C_{\ell-1}$ possibilities. By similar logic, there are $\ell-2$ ways to obtain a $q^{\ell}$ charge of the form $(()\dots)^{\ell}$, where $\dots$ is a path of length $\ell-2$. To obtain the expression for $p_{\rm in}(\ell)$, we consider both words of the form $(()\dots)^{\ell}$ and $(\dots())^{\ell}$, of which there are $2~{}C_{\ell-2}$; however, care must be taken to not count configurations like $(()\dots())^{\ell}$ twice. Since $\dots$ in these configurations must be a Dyck path of semilength $\ell-3$, the expression for $p_{\rm in}(\ell)$ is given by the following:
@@
The expressions in eqs. 15 and 16 for $p_{\rm in}(\ell)$ and $p_{\rm out}(\ell,L)$ are conditioned on finding $q^{\ell}$ on a particular link; however, in order to obtain the expression for $\langle q_{0}j_{n}\rangle$, these conditional probabilities need to be multiplied by the probability to have matched parentheses $q_{0}^{\ell}$ in the first place. To produce the relevant expressions, we write down the probability $p_{q}(\ell,L)$ of finding a $q^{\ell}$ charge on any particular link in a system of $L$ sites. Overall, there are $\binom{L}{L/2}$ possible states that the system can be found in as a while. Since there are $C_{\ell-1}$ possible Dyck paths corresponding to a $q^{\ell}$ charge, and since there are $\binom{L-2\ell}{(L-2\ell)/2}$ possible states that the the $L-2\ell$ sites outside of $q^{\ell}$ may take, $p_{q}(\ell,L)$ is given by the following expression:
@@ Appendix C Numerical results for the Fredkin matched parentheses charge
Figure 3: Numerical results (a) The dynamical exponent $z$ extracted from structure factor data for the Fredkin Markov process. At long times, $z$ appears to saturate to a value slightly larger that 5/2. The structure factor collapses for the original $U(1)$ charge (b) and the matched parentheses charge (c), using $z=5/2$, shown for times $t=200-400$.
@@ D.2 The matched parentheses charge in Motzkin
Taking into account all of the considerations outlined above, the Motzkin Hamiltonian may be rewritten in terms of the matched parentheses charge. Writing the hard-core bosonic ladder operators for matched parentheses charges of length $\ell$ whose midpoint position is $i$ as $a_{i}^{\ell({\dagger})}$, the Hamiltonian is given by:
@@ D.3.1 Power-law tails
The main text outlines how all static correlations in spin-$\frac{1}{2}$ systems may be traced back to the fact that each endpoint of an irreducible word uniquely belongs to a single irreducible word. The exact same argument holds for spin-1 systems. In a spin-1 system, the existence of a single $q_{0}^{1}$ charge precludes the simultaneous existence of any $q_{\pm\ell}^{\ell}$ charge; the only differences between the integer and half-integer spin models is that here, $q_{0}^{1}$ is a flat and $q_{\pm\ell}^{\ell}$ may live on either a link or a site.
@@ D.4.2 Charge-current correlations
The origin of charge-current correlations may be understood in direct analogy to the source of charge-current correlations in Fredkin, as outlined in B.2. In Motzkin, the movement of all $q^{\ell}$ charges, for $\ell\geq 1$, is facilitated by a single $q^{1}$ charge passing through one of the endpoints of the $q^{\ell}$ charge. As the $q^{1}$ charge disrupts one of the endpoints of the $q^{\ell}$ charge, the larger charge changes both length and position to become a $q_{i\pm 1}^{\ell\pm 1}$ charge.
As already noted, the continuity equation contains contributions from two different types of currents of the matched parentheses charges ($j$ and $j^{\prime}$) due to differences between the movement of flats ($q^{1}$ charges) and charges associated with $\ell\geq 2$ Motzkin words. During the process of passing through the endpoint of a larger word, the $q^{1}$ charge moves two, since the $q^{1}$ charge is may only live on sites. In contrast, the other charge involved in this process will move by only one space as it increases or decreases in length by one. Therefore, since $\ell=1$ charges move different amounts than $\ell\geq 2$ charges, the continuity equation associated with the movement of $q_{n}$ cares whether the matched parentheses current at $n$ is caused by a $\ell=1$ or $\ell\geq 1$ charges.
@@ D.4.4 Motzkin spin-wave energy $E_{\rm SW}(k)$
The energy of the spin-wave state splits into two parts, $E_{\rm SW}=E_{\rm SW}^{0}(k)+E_{\rm SW}^{1}(k)$, as shown in 30. Each $E_{\rm SW}^{m}(k)$ has contributions from both $j_{n}$ currents and $j^{\prime}_{n}$ currents:

### v5

@@ B.2 Source of charge-current correlations
In the picture of the matched parentheses charge, it becomes clear that the source of all Fredkin dynamics is the movement of charges of semilength $\ell=1$, which are matched parentheses of the form (). When an () charge moves, it necessarily disrupts the end parenthesis of another charge of semilength $\ell$. This can be pictured as the () moving past one of the matched end parentheses of the charge of semilength $\ell$, which we will call moving “inside” or “outside” of the charge of semilength $\ell$. The movement of the () charge triggers the process in which the charge of semilength $\ell$ hops by a single site to link $n=\pm 1$ while changing semilength by one to ±$\ell\pm 1$.
@@
Since the only source of dynamics from the Fredkin Hamiltonian involve locally moving () charges, all nontrivial correlations between a charge sitting on link 0 and a current on site $n$ must come from the coincidence of two scenarios: first, that the charge sitting on link 0 corresponds to matched parentheses separated by ±$\ell\pm 1$, and second, that a () charge on link $n=\pm(\ell\mp 1)$ disturbs the end parenthesis of the charge on link 0. This statement has an intuitive interpretation– a current may only be correlated with a charge of semilength $\ell$ if a () charge either moves inside or outside the charge, which consequently changes the separation of the charge to ±$\ell\pm 1$. In order to move inside or outside of the charge of semilength $\ell$ at a given timestep, since there are no non-local () hops encoded in the Fredkin Hamiltonian, there must be a () charge located directly adjacent to one of the matched end parentheses of the charge of semilength $\ell$ immediately before the hop. Therefore, the task of understanding the static charge-current correlations in Fredkin reduces to finding the probability of these configurations. Exact expressions for these probabilities are derived in the following section.
@@ B.3 Probabilities
First, we focus on the process $()(\dots)^{\ell}\rightarrow(()\dots)^{\ell+1}$, in which $(\dots)^{\ell}$ pictorially represents some matched parentheses charge of semilenght $\ell$. We denote the probability of finding a configuration of the form $()(\dots)^{\ell}$ or $(\dots)^{\ell}()$ conditional on having a $(\dots)^{\ell}$ as $p_{\rm out}(\ell)$. From direct counting, we find that $p_{\rm out}(\ell)=7/16$ in the thermodynamic limit since the system is locally at infinite temperature. For a system of $L$ sites in the $S^{z}_{\rm tot}=0$ ensemble, however, the expression for $p_{\rm out}(\ell,L)$ becomes dependent on both $\ell$ and $L$ due to additional correlations that arise from finite-size effects.
Since any Dyck word contains equal numbers of up and down spins, the $L-2\ell$ sites in the system that are not part of the $(\dots)_{\ell}$ charge must also contain the same number of up and down spins when the system is at $S^{z}_{\rm tot}=0$. There are $\binom{L-2\ell}{(L-2\ell)/2}$ possible states for these $L-2\ell$ sites. Consider a Dyck word of semilength $\ell+1$ of the form $()(\dots)^{\ell}$; since $()(\dots)^{\ell}$ also contains only matched sets of parentheses, the $L-(\ell+2)$ sites in the system that are not part of $q^{\ell}$ must contain the same number of up and down spins. Therefore, there are $\binom{L-2\ell-2}{(L-2\ell-2)/2}$ possible configurations for these $L-2\ell-2$ sites. Finally, taking care to avoid double-counting the $\binom{L-2\ell-4}{(L-2\ell-4)/2}$ configurations of the form $()(\dots)^{\ell}()$, we obtain the following expression for $p_{\rm out}(\ell,L)$:
@@
Next, we focus on the process $(()\dots)^{\ell}\rightarrow()(\dots)^{\ell-1}$. We denote the probability of finding a configuration of the form $(()\dots)^{\ell}$ conditional on having a $(\dots)^{\ell}$ as $p_{\rm in}(\ell)$. Unlike when a charge of semilength $\ell=1$ sits directly outside of a charge of semilength $ell$, a charge of semilength $\ell=1$ sitting directly inside of a $q^{\ell}$ charge is subject to strong correlations from being embedded within a Dyck word. In total, there are $C_{\ell-1}$ possible Dyck words of semilength $\ell$, where $C_{\ell}$ are Catalan numbers. Recall that this was obtained from there fact that of the $C_{\ell}$ Dyck word of semilength $\ell$, only those of the form $(\dots)^{\ell}$ correspond to matched parentheses charges; therefore, the $\dots$ in $(\dots)^{\ell}$ must also be a Dyck word of semilength $\ell-1$, of which there are $C_{\ell-1}$ possibilities. By similar logic, there are $\ell-2$ ways to obtain a charge of the form $(()\dots)^{\ell}$, where $\dots$ is a path of length $\ell-2$. To obtain the expression for $p_{\rm in}(\ell)$, we consider both words of the form $(()\dots)^{\ell}$ and $(\dots())^{\ell}$, of which there are $2~{}C_{\ell-2}$; however, care must be taken to not count configurations like $(()\dots())^{\ell}$ twice. Since $\dots$ in these configurations must be a Dyck path of semilength $\ell-3$, the expression for $p_{\rm in}(\ell)$ is given by the following:
@@
The expressions in eqs. 17 and 18 for $p_{\rm in}(\ell)$ and $p_{\rm out}(\ell,L)$ are conditioned on finding a charge of semilength $\ell$ on a particular link; however, in order to obtain the expression for $\langle q_{0}j_{n}\rangle$, these conditional probabilities need to be multiplied by the probability to have a matched parentheses charge of semilength $\ell$ on link $n=0$ in the first place. To produce the relevant expressions, we write down the probability $p_{q}(\ell,L)$ of finding a charge of semilength $\ell$ on any particular link in a system of $L$ sites. Overall, there are $\binom{L}{L/2}$ possible states that the system can be found in as a while. Since there are $C_{\ell-1}$ possible Dyck paths corresponding to a charge of semilength $\ell$, and since there are $\binom{L-2\ell}{(L-2\ell)/2}$ possible states that the the $L-2\ell$ sites outside of $(\dots)_{\ell}$ may take, $p_{q}(\ell,L)$ is given by the following expression:
@@ Appendix C Numerical results for the Fredkin matched parentheses charge
Figure 3: Numerical results (a) The dynamical exponent $z$ extracted from structure factor data for the Fredkin Markov process. At long times, $z$ appears to saturate to a value slightly larger that 5/2, although we cannot rule out logarithmic corrections. The structure factor collapses for the original $U(1)$ charge (b) and the matched parentheses charge (c), using $z=5/2$, shown for times $t=200-400$. For all numerics, system sizes were chosen such that finite-size effects are negligible on the timescales shown.
@@ D.2 The matched parentheses charge in Motzkin
Taking into account all of the considerations outlined above, the Motzkin Hamiltonian may be rewritten in terms of the matched parentheses charge. Writing the hard-core bosonic ladder operators for matched parentheses charges of length $\ell$ whose midpoint position is $i$ as $a_{i}^{\ell({\dagger})}$, the Hamiltonian in the $S^{z}_{\rm tot}$ charge sector with PBCs is given by:
@@ D.3.1 Power-law tails
The main text outlines how all static correlations in spin-$\frac{1}{2}$ systems may be traced back to the fact that each endpoint of an irreducible word uniquely belongs to a single irreducible word. The exact same argument holds for spin-1 systems. In a spin-1 system, the existence of a single flat on site 0 precludes the simultaneous existence of any charge of length $\ell$ on either a link or site $\pm\ell$ away from the flat.
@@ D.4.2 Charge-current correlations
The origin of charge-current correlations may be understood in direct analogy to the source of charge-current correlations in Fredkin, as outlined in B.2. In Motzkin, the movement of any charge of length $\ell>2$ with position $i$ is facilitated by a single flat passing through one of its endpoints. As the flat disrupts one of the endpoints of the larger charge, the larger charge changes in both length to ±$\ell\pm 1$ and position to ${i\pm 1}$.
As already noted, the continuity equation contains contributions from two different types of currents of the matched parentheses charges ($j$ and $j^{\prime}$) due to differences between the movement of flats and charges with length $\ell\geq 2$. During the process of passing through the endpoint of a larger word, the flat changes position by two, since the flat may only live on sites. In contrast, the other charge involved in this process will move by only one space as it increases or decreases in length by one. Therefore, since $\ell=1$ charges (flats) move different amounts than $\ell\geq 2$ charges, the continuity equation associated with the movement of $q_{n}$ cares whether the matched parentheses current at $n$ is caused by a $\ell=1$ or $\ell\geq 1$ charges.
@@ D.4.4 Motzkin spin-wave energy $E_{\rm SW}(k)$
The energy of the spin-wave state splits into two parts, $E_{\rm SW}=E_{\rm SW}^{0}(k)+E_{\rm SW}^{1}(k)$, as shown in 32. Each $E_{\rm SW}^{m}(k)$ has contributions from both $j_{n}$ currents and $j^{\prime}_{n}$ currents: